OpenStudy (anonymous):

A tub contains two faucets. Faucet A can fill the tub in 15 minutes and faucet B can fill the tub in 10 minutes. How long will it take to fill the tub using both faucets? (A) 6 min (B) 7.5 min (C) 8 min (D) 25 min (E) Not enough info

hero (hero):

If: A = amount of time it takes faucet A to fill the tub alone B = amount of time it takes faucet B to fill the tub alone t = amount of time it takes both faucets to fill the tub Then the following equation applies: \[\frac{AB}{A + B} = t\] In this case: A = 15 B = 10 so: \[\frac{(15)(10)}{15 + 10} = t\] Solve for \(t\)

OpenStudy (anonymous):

@Hero I got 6 c:

hero (hero):

Correct

hero (hero):

Working together, Faucets A and B will fill the tub in 6 minutes.

OpenStudy (anonymous):

Thank you! c: